find the values for me and b in the following system so that the solution to the system is (-3,4)
5x+7y=b
mx+y=22

Answer by venugopalramana(3286)

find the values for me and b in the following system so that the solution to the system is (-3,4)
5x+7y=b.........5*(-3)+7*4=b=13
mx+y=22.........m*(-3)+4=22.....-3m=22-4=18...m=-18/3=-6